Part A Standardized stock price indicators in three different countries over a week are listed below....

Part A

Standardized stock price indicators in three

different countries over a week are listed below. An analyst is interested in knowing if

the stock markets of these different countries are dependent on one another. The

data set and a partial ANOVA table for this study are provided below.

I

II

III

890

900

905

899

900

900

900

887

896

905

906

928

871

893

899

910

900

934

Source of variation     

SS

DF

MS

F

Treatment                   

748

2

374

???

Error                            

2526

???

???

            

Total                            

3274

???

            

Compute the MSE and the F statistic

  1. MSE = 374. F = 2.22
  2. MSE = 168.4. F = 2.22
  3. MSE = 2.22. F = 15
  4. MSE = 2,526. F = 2.22
  5. None of the above

Part B

Suppose the p-value for the test is 0.143. At the 0.05 level of significance, how do you conclude?

  1. Do not reject H0. There is no evidence that the means are significantly different
  2. Do not reject H0. Evidence exists that the means are significantly different
  3. Reject H0. There is no evidence that the means are different
  4. Reject H0. P-value is greater than alpha
  5. None of the above statements is correct

Part C

Calculate the Tukey Criterion (T) for use in a Tukey pairwise comparisons test. Use alpha of 0.05.

  1. T = 3.67
  2. T = 19.05
  3. T = 19.44
  4. None of the above

Part D

Which of the pairs of sample means are statistically significant?

  1. Sample means for Groups 1 and 2
  2. Sample means for Groups 1 and 3
  3. Sample means for Groups 2 and 3
  4. None is significant
Answer #1

The number of columns, k = 3

The Total observations, N = 18

DF Treatment = n - 1 = 2

DF Error = N - k = 18 - 3 = 15

DF Total = DF treatment + DF error = 2 + 15 = 17

MSTR = SS treatment/DF treatment = 748/2 = 374

__________________________________________

PART A: Option b

MSE = SS Error/DF error = 2526/15 = 168.4

F statistic = MSTR/MSE = 374/168.4 = 2.22

__________________________________________

PART B: Option a

If p value is < alpha, then Reject H0

Since p value (0.143) is > alpha (0.05), Do not Reject H0. There is no evidence that the means are significantly different.

_________________________________________

PART C: Option a

Tukeys Crirerior for pairwise comparison at alpha = 0.05

For k = 3 and DF error = 15, the Tukey Value T (from tables) = 3.67

__________________________________________

Part D: Option d, none is significant.

It is given by

HSD = rac{M_i-M_j}{sqrt{rac{MS_{error}}{n}}}

where Mi and Mj are the 2 means being compared and their positive difference is taken. n = number of replicates in each sample. Here n = 6, MSerror 168.4

/168.4 eTTo

The Rule is that if Tukeys Observed is > Tukeys Critical, then there is a significant difference between the groups.

M1 = 895.83, M2 = 897.67 and M3 = 910.33

Group 1 and Group 2 : M1 - M2 = Absolute (895.83 - 897.67) = 1.84

Tukeys Observed = 1.84/5.28 = 0.348.

Since Tukeys observed (0.348) is < Tukeys Critical (3.67), there isn't a significant difference between the 2 groups.

Group 1 and Group 3 : M1 - M3 = Absolute (895.83 - 910.33) = 14.5

Tukeys Observed = 14.5/5.28 = 2.746.

Since Tukeys observed (2.746) is < Tukeys Critical (3.67), there isn't a significant difference between the 2 groups.

Group 2 and Group 3 : M2 - M3 = Absolute (897.67 - 910.33) = 12.66

Tukeys Observed = 12.66/5.28 = 2.398.

Since Tukeys observed (2.398) is < Tukeys Critical (3.67), there isn't a significant difference between the 2 groups.

__________________________________________________________________